A group of students are on a tour. The total fare is $120 and this is to be shared equally among the students. If two more students join the tour, each will pay two dollars less. Find the original number of students in the group.
Answer by checkley77(12569)

120/X=120/(X+2)+2
120/X=(120+2[X+2])/X+2
120(X+2)=X(120+2X+4)
120X+240=120X+2X^2+4X
2X^2+120X-120X+4X-240=0
2X^2+4X-240=0
2(X^2+2X-120)=0
2(X-10)(X+12)=0
X-10=0
X=10 ANSWER FOR THE ORIGINAL NUMBER OF STUDENTS.
10+212 THE FINAL NUMBER OF STUDENTS.
PROOF
120/10=120/12+2
12=10+2
12=12